OpenStudy (lovelyharmonics):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i

OpenStudy (anonymous):

If 2 + 5 i is a root, them 2-5i is also. (x−4)(x−(2+5i))(x−(2−5i))(x+8) Do you know how to expand it?

OpenStudy (anonymous):

ok the easy part is to start with \[(x-4)(x+14)\] what will give you a polynomial whose zeros are \(4\) and \(-14\) the harder part is coming up with a quadratic whose zeros are \(5+8i\) and \(5-8i\) there are several ways to do it i can show you an easy one, and a real real easy one if you like

OpenStudy (lovelyharmonics):

can you not create a foil problem like you did with the first?

OpenStudy (anonymous):

one way is to work backwards put \[x=5+8i\] subtract \(5\) get \[x-5=8i\] square both sides carefully and get \[(x-5)^2=(8i)^2\] or \[x^2-10x+25=-64\] then add \(64\) to get \[x^2-10x+89\] as your quadratic

OpenStudy (anonymous):

there is an even snappier way, but it requires memorizing something the quadratic with zeros \(a+bi\) and\(a-bi\) is \[x^2-2ax+(a^2+b^2)\] in your case \(5+8i\) you have \(a=5,b=8\) so it is \[x^2-2\times 5x+(5^2+8^2)\] or \[x^2-10x+89\] as before

OpenStudy (anonymous):

final job is to multiply out \[(x-4)(x+14)(x^2-10x+89\] if it was me, i would cheat

OpenStudy (lovelyharmonics):

hold on where did you just get (or x2−10x+25=−64 then add 64 to get x2−10x+89 as your quadratic) this part from?

OpenStudy (anonymous):

lets take it step by step starting with \[x=5+8i\] which is what you were given

OpenStudy (anonymous):

when you subtract 5 you get \[x-5=8i\] so far so good?

OpenStudy (lovelyharmonics):

yes c:

OpenStudy (anonymous):

now square both sides

OpenStudy (lovelyharmonics):

x^2+25=64(-1) =-64

OpenStudy (anonymous):

on the right, you get \((8i)^2=8i\times 8i=8^2i^2=64(-1)=-64\)

OpenStudy (anonymous):

but you have to square carefully on the left

OpenStudy (anonymous):

\[(x-5)^2=(x-5)(x-5)=x^2-5x-5x+25=x^2-10x+25\]

OpenStudy (lovelyharmonics):

oh okay you foil the left

OpenStudy (anonymous):

i call it multiplying but i suppose you could say ...

OpenStudy (anonymous):

in any case now you see why the left is \[x^2-10x+25=-64\] and once you add \(64\) you get \[x^2-10x+89\]

OpenStudy (lovelyharmonics):

okay and then what

OpenStudy (anonymous):

you gotta multiply out this ugly mess \[(x-4)(x+14)(x^2-10x+89)\] like i said, i would cheat want to check it?

OpenStudy (lovelyharmonics):

one sec im working it our c:

OpenStudy (anonymous):

k let me know when you want the answer

OpenStudy (lovelyharmonics):

k so is it x^4-30x^3-487x^2+1450x-4984

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%28x-4%29%28x%2B14%29%28x^2-10x%2B89%29 looks like it is \(x^4-67 x^2+1450 x-4984\)

OpenStudy (anonymous):

hard to keep it all straight, that is why i said i would cheat almost impossible not to make some kind of mistake

OpenStudy (lovelyharmonics):

okay thanks

OpenStudy (anonymous):

yw use wolfram, it is your friend